# WinkTPG: An Execution Framework for Multi-Agent Path Finding Using Temporal Reasoning

### v1

@@ 8 Conclusion
We introduced WinkTPG, an execution framework for efficiently and robustly executing MAPF plans in environments with execution uncertainty. WinkTPG uses our proposed kinodynamic planning method kTPG to generate collision-free speed profiles for all agents. It incorporates location updates from agents to replan dynamically while ensuring robust execution. We evaluated our methods in the MAPF domain, showing up to 51.7% improvement in solution quality compared to existing methods, while generating speed profiles for up to 1,000 agents in 1 second. Future work includes exploring the integration of WinkTPG with advanced MAPF replanning techniques such as rolling-horizon conflict resolution (Li et al. 2021) or switchable TPG (Feng et al. 2024).

### v2

@@ VIII Conclusion
In this paper, we introduced WinkTPG, an execution framework for efficiently and robustly executing MAPF plans in environments with execution timing uncertainty. WinkTPG uses our proposed kinodynamic planning method kTPG to generate collision-free speed profiles for all agents. It incorporates location updates from agents to replan dynamically, while providing deterministic guarantees under bounded delays and probabilistic guarantees under stochastic models. Empirically, we evaluated our methods in the MAPF domain, showing up to 51.7% improvement in solution quality compared to existing methods, while generating speed profiles for up to 1,000 agents in 1 second. Under multiple uncertainty models, WinkTPG achieved the intended robustness guarantees in practice. The results also showed that WinkTPG outperforms ADG under moderate uncertainty. We further validated the practical applicability of WinkTPG through experiments in a high-fidelity physics-based simulator and real-world mobile robots, confirming that the generated speed profiles are physically executable and robust to execution variability. Future work includes exploring the integration of WinkTPG with advanced MAPF replanning techniques such as rolling-horizon conflict resolution [15] or switchable TPG [8]. Another important direction is to extend the uncertainty analysis beyond the current setting. This includes heavy-tailed disturbances, correlated delays arising from shared infrastructure, and adversarial failures, none of which are explicitly modeled in the present framework. Finally, extending WinkTPG to incorporate decentralized planning and execution mechanisms could further enhance its scalability and robustness in highly dynamic and large-scale scenarios.